OpenStudy (sambam):

In a solution with a pH of 4, the [OH-] is?

OpenStudy (frostbite):

I'm gonna write some theory then you do the calculation: Water protolytic properties make the following true for pure water and dilute aqueous solutions at 25 ° C: \[[H _{3}O ^{+}]=[OH ^{-}]=1,0 * 10^{-7} M <-> [H _{3}O ^{+}]*[OH ^{-}]=1,0*10^{-14} M\] pH is now defined as negative logarithm of the concentration of H 3 O +, while pOH is defined as the negative logarithm of the OH - concentration: \[pH=-\log([H ^{+}])\] \[pOH=-\log([OH ^{-}])\] We take the logarithm on both sides in the first written equation: \[1,0*10^{-14} M = K _{w}=[H _{3}O ^{+}]*[OH ^{-}] <-> \log(K _{w})=pH*pOH\] In other terms there following most apply at 25 ° C: \[pH*pOH=14,0\]

OpenStudy (frostbite):

You know where to begin or need further help?

OpenStudy (frostbite):

I give you some hints. Start by using the last equation pH*pOH=14,0 to find pOH Then use the pOH=-log([OH-]) equation... solved for [OH-] it is going to look like this: \[[OH ^{-}]=10^{-pOH}\]

OpenStudy (frostbite):

sorry meant: \[pH+pOH=14,0\]